# Suiren-1.0 Technical Report: A Family of Molecular Foundation Models

### 13 Apr 2026 v2

@@ 3.3 Pre-training Evaluation
We sample a validation subset from Qo2mol containing more than 1M conformers across different molecular scales. This subset provides a stable benchmark for monitoring pre-training quality. As shown in Table 1, compared with baseline methods, Suiren-Base achieves highly accurate results on both energy and force prediction. Optimized-trajectory endpoint structure and endpoint energy are substantially more challenging targets, yet Suiren-1.0 still attains strong accuracy.
Suiren-0.0 is an internal transitional model trained with less compute and a weakened training recipe. Because it uses exactly the same training set as Suiren-Base, its results directly reflect the gains brought by the improved training strategy and algorithmic refinements in Suiren-1.0. EquiformerV2 [liao2023equiformerv2] and eSCN [passaro2023reducing] are strong backbones; due to compute constraints, we train them on a 20M Qo2mol subset. Their results further support the effectiveness of the Suiren-1.0 training pipeline.
We also compare Atomic Energy MAE and Force MAE with UMA-family models on the organic benchmark set (OMol). Our energy prediction is comparable or better ($0.258$ vs. >0.33>0.33), while our force prediction shows a much larger improvement ($0.510$ vs. >2.90>2.90). Note that this UMA comparison is intended only to provide a rough performance reference, since the training and evaluation datasets are not identical across methods.
Finally, we evaluate the continued pre-training model, Suiren-Dimer. Compared with intra-molecular settings, inter-molecular trajectories are more complex. Accordingly, Suiren-Dimer is weaker than Suiren-Base on endpoint structure and endpoint energy prediction, but it still maintains strong performance on energy and force prediction.
Table 1: Comparison among energy/forces prediction models. The best results are shown in bold.
EquiformerV2 eSCN Suiren-0.0 Suiren-Base (1.0) Suiren-Dimer (1.0) # Total Params 31M 168M 1.5B 1.8B 1.8B Molecular Energy MAE (meV) $\downarrow$ 206.15 268.1 10.82 9.08 10.30 Atomic Energy MAE (meV) $\downarrow$ 5.854 7.613 0.307 0.258 0.292 Atomic Forces MAE (meV / Å) $\downarrow$ 29.3 37.6 0.979 0.510 1.294 Optimized Structure MAE (Å) $\downarrow$ - - 0.02895 0.02467 0.06830 Optimized Energy MAE (meV) $\downarrow$ - - 28.187 16.886 58.325
@@ 5.1.1 MoleHB
We introduce Molelecular handbook (MoleHB), a comprehensive molecular property prediction benchmark encompassing 40+ heterogeneous tasks. The benchmark spans several critical scientific domains, including safety, structural, critical and saturation, energetic, thermal, solution, transport and fluctuation properties. All data points are inferred from canonical handbook [yaws1999chemical] and have been rigorously validated via wet-lab experiments to ensure high-fidelity, stable values. We propose two evaluation protocols: (1) Random split: a standard random split to evaluate performance under similar data distributions; and (2) Scaffold split: a strategy where molecules with larger atom counts are assigned to the validation set to assess the model’s structural extrapolation capabilities. To facilitate reproducible research, we provide the MoleHB dataset and two splitting protocols on modelscope.cn/datasets/ajy112/MoleHb
The main paper primarily considers results from the Random split. Results of Scaffold split are shown in Appendix C.
@@
Train/validation split ratio (Random and Scaffold) 0.8:0.2
@@ Appendix B The explanation of evaluation metrics in Figure 1 and Figure 2
Appendix C Evaluation of MoleHB Scaffold split
To further explore the generalization capability of foundation models under distribution shift, we systematically evaluated MoleBERT, Uni-Mol v1, Uni-Mol v2, and Suiren-ConfAvg on the scaffold-split subset of MoleHB, a setting that rigorously tests extrapolation to unseen molecular scaffolds. The comprehensive results across eight property categories are summarized in Tables 1522.
@@ Appendix C Evaluation of MoleHB Scaffold split
As anticipated, scaffold splitting induces a non-trivial distribution shift, leading to performance degradation across all methods relative to random-split evaluations. Nevertheless, Suiren-ConfAvg demonstrates superior robustness: it achieves the lowest mean absolute error (MAE) on properties across all categories, with relative improvements ranging from 4.68% (critical pressure) to 91.20% (Helmholtz energy of formation) over the strongest baseline. Notably these confidence-weighted ensemble approaches achieve substantial improvements particularly on energetic and critical properties, underscoring the effectiveness of confidence-weighted ensemble strategies in mitigating scaffold-induced generalization gaps.
@@
  • Thermal and fluctuation properties (Tables 1917): The density of liquid shows moderate performance variation across splits (Suiren-ConfAvg: 0.00851), consistent with its dependence on atomic composition, with a 29.67% improvement over the best baseline. In contrast, heat capacities exhibit substantial scaffold sensitivity, with Suiren-ConfAvg achieving 41.50%–88.63% improvements for liquid and solid heat capacities respectively, while showing 72.33% improvement for gas heat capacity.
@@
Statistical considerations and limitations
While the scaffold split provides a rigorous test of extrapolation, we acknowledge that the reported MAE values are point estimates without confidence intervals. Future evaluations should incorporate bootstrap resampling or cross-validation over multiple scaffold partitions to assess result stability. Additionally, the observed performance gaps may partially reflect differences in model capacity and pre-training data scale.
@@
The pronounced robustness of Suiren-ConfAvg on energetically complex and scaffold-sensitive properties suggests that integrating physics-aware pre-training objectives with uncertainty-aware inference mechanisms can substantially improve out-of-distribution generalization.

### 28 Apr 2026 v4

@@ 3.3 Pre-training Evaluation
We sample validation subsets from Qo2mol containing more than 1M conformers across different molecular scales. Suiren-Base and Suiren-Dimer use different subsets. As shown in Table 1, Suiren-Base achieves highly accurate results on both energy and force prediction. Optimized-trajectory endpoint structure and endpoint energy are substantially more challenging targets, yet Suiren-1.0 still attains strong accuracy. Note that Suiren-0.0 is an internal transitional model trained with less parameters and a weakened backbone architecture. It uses the same training set and hyperparameters as Suiren-Base (1.0). We attribute Suiren-Base’s improvements to increased parameters, optimized backbone architecture, and refined training algorithms. Additionally, we evaluate the continued pre-training model, Suiren-Dimer. Compared with intra-molecular settings, inter-molecular trajectories are more complex. Accordingly, Suiren-Dimer is weaker than Suiren-Base on endpoint structure and endpoint energy prediction, but it still maintains strong performance on energy and force prediction. The comparison between Suiren-Base and Suiren-Dimer is indirect, as they used different validation sets.
Table 1: Comparison among energy/forces prediction models.
Suiren-0.0 Suiren-Base (1.0) Suiren-Dimer (1.0) # Total Params 1.5B 1.8B 1.8B Molecular Energy MAE (meV) $\downarrow$ 10.82 9.08 10.30 Atomic Energy MAE (meV) $\downarrow$ 0.307 0.258 0.292 Atomic Forces MAE (meV / Å) $\downarrow$ 1.472 0.770 1.294 Optimized Structure MAE (Å) $\downarrow$ 0.02895 0.02467 0.06830 Optimized Energy MAE (meV) $\downarrow$ 28.187 16.886 58.325
@@ 5.1.1 MoleHB
We introduce Molelecular handbook (MoleHB), a comprehensive molecular property prediction benchmark encompassing 40+ heterogeneous tasks. The benchmark spans several critical scientific domains, including safety, structural, critical and saturation, energetic, thermal, solution, transport and fluctuation properties. All data points are inferred from canonical handbook [yaws1999chemical] and have been rigorously validated via wet-lab experiments to ensure high-fidelity, stable values. We propose two evaluation protocols: (1) Random split: a standard random split to evaluate performance under similar data distributions; and (2) Size-Stratified split: a strategy where molecules with larger atom counts are assigned to the validation set to assess the model’s structural extrapolation capabilities. To facilitate reproducible research, we provide the MoleHB dataset and two splitting protocols on https://modelscope.cn/datasets/ajy112/MoleHB
The main paper primarily considers results from the Random split. Results of Size-Stratified split are shown in Appendix C.
@@
Train/validation split ratio (Random and Size-Stratified) 0.8:0.2
@@ Appendix B The explanation of evaluation metrics in Figure 1 and Figure 2
Appendix C Evaluation of MoleHB Size-Stratified split
To further explore the generalization capability of foundation models under distribution shift, we systematically evaluated MoleBERT, Uni-Mol v1, Uni-Mol v2, and Suiren-ConfAvg on the Size-Stratified split subset of MoleHB, a setting that rigorously tests extrapolation to unseen molecular sizes. The comprehensive results across eight property categories are summarized in Tables 1522.
@@ Appendix C Evaluation of MoleHB Size-Stratified split
As anticipated, Size-Stratified split induces a non-trivial distribution shift, leading to performance degradation across all methods relative to random-split evaluations. Nevertheless, Suiren-ConfAvg demonstrates superior robustness: it achieves the lowest mean absolute error (MAE) on properties across all categories, with relative improvements ranging from 4.68% (critical pressure) to 91.20% (Helmholtz energy of formation) over the strongest baseline. Notably these confidence-weighted ensemble approaches achieve substantial improvements particularly on energetic and critical properties, underscoring the effectiveness of confidence-weighted ensemble strategies in mitigating size-induced generalization gaps.
@@
  • Thermal and fluctuation properties (Tables 1917): The density of liquid shows moderate performance variation across splits (Suiren-ConfAvg: 0.00851), consistent with its dependence on atomic composition, with a 29.67% improvement over the best baseline. In contrast, heat capacities exhibit substantial size sensitivity, with Suiren-ConfAvg achieving 41.50%–88.63% improvements for liquid and solid heat capacities respectively, while showing 72.33% improvement for gas heat capacity.
@@
The pronounced robustness of Suiren-ConfAvg on energetically complex and size-sensitive properties suggests that integrating physics-aware pre-training objectives with uncertainty-aware inference mechanisms can substantially improve out-of-distribution generalization.